OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

ok lets say you have 20x = 30 the answer to the equation would be x = 30/20 or 3/2

OpenStudy (anonymous):

your goal was to find the value of x by doing this you have solve the unknown in the equation

OpenStudy (anonymous):

A solution to an equation means that it satisfies the equal sign and evaluates as true. For example: 2x-5=12 --> x=(12+5)/2 --> x=8.5 You can check this: 2*8.5-5=12 <-- this equality is satisfied and evaluates as true. 12=12 Twelve is equal to twelve; how very true!